# GTransPDM: A Graph-embedded Transformer with Positional Decoupling for Pedestrian Crossing Intention Prediction

### v1

@@ IV-A Datasets
The data used in the PIE experiments was split into train, validation (val), and test at a ratio of 0.5:0.1:0.4, with random splitting based on person IDs. For comparison with existing methods (Table I), the performance following the data split in [3] was tested, with 4770, 1332, and 3816 samples in train, val, and test sets respectively. For the JAAD dataset, the JAAD_all subset and adhered to the default data split provided by the benchmark were used, including 8613, 1265, and 6732 sequences for each part. Furthermore, 0.5 second ($T=16$) observations were utilized to predict pedestrian crossing intention after 1-2 seconds ($TTE\in[30,60]$), sampled data based on the sliding window with an overlap of 0.6, 0.8 for PIE and JAAD dataset, respectively.
IV-B Implementation Details
@@
In the network architecture, the dimension of three encoder layers $C_{d}$, hidden layers in GCN $d_{hid}$, and $d_{m}$ in Transformers are all set to 64. With the concatenate operation in the feature fusion block, the channel dimension of the fused feature is $1923\times C_{d}=$ 192. In the transformer encoder, $dropout=0.1$, 4 attention heads, and 4 stacked layers were applied to extract temporal information. The reference lines in PDM were set to one line between points: $A=(0,H)$ and $B=(W/2,Y_{min})$ while another between: $B=(W/2,Y_{min})$ and $C=(W,H)$, as shown in Fig. 2 (a), where $Y_{min}$ is the minimum $y_{t}$ of all samples in each dataset. The minor differences in coordinates have little impact on model performance in subsequent experiments, as shown in Table II. The keypoints were offline generated by the pre-trained Alphapose [42], while the additional neck, hip, and body center points were further calculated based on the average of adjacent points.
@@
IV-C Evaluations
The effectiveness of our method was evaluated using Accuracy, Area Under Curve (AUC), F1 Score, Precision, and Recall, which are commonly used in existing studies. Table I compares our method with recent approaches from the past five years, including Transformer-based IntFormer [10] and PIT [24], the graph-based Pedestrian Graph+[45], and several RNN-based methods. Except BiPed[44] and Pedestrian Graph+ [45], other solutions show the same data splits and configurations as ours, following [3]. Additionally, we trained and tested MultiRNN[46], SingleRNN[17], and SFRNN[13], published before the benchmark released, using the codebase from [3]. While PIT [24] captures traffic agent interactions using a ViT-based  [23] network with 16 stacked Transformer layers, incurring high computational costs, our lightweight network achieves superior performance, reaching 0.92 accuracy and 0.86 F1 in PIE using only position and ego-vehicle motion encoders, with 0.86 precision when incorporating the pose encoder. However, as shown in Table I, most PIE metrics are higher without $X_{ke}$, likely due to the benchmark’s video-based data split, reducing the randomness of pedestrian crossing patterns. To ensure a more balanced comparison, we split the data by pedestrian IDs in later experiments. Regarding JAAD, we achieves 0.87 accuracy and 0.64 precision, surpassing most existing methods.
Fig. 3 shows the performance based on different TTE settings, ranging from 0 to 60 frames. A smaller TTE value indicates a shorter prediction interval. In the PIE experiment, we observed that as TTE decreased, the model’s performance improved rapidly. Without a TTE interval (TTE = 0), the accuracy and F1 score reached 99.32% and 98.73%, respectively. However, in the JAAD experiments, the performance improved only when very close to the action start time, mostly because the JAAD_all subset contains a large number of pedestrians who are far away and do not interact with the driver, which confuses the model’s ability to identify crossing behavior that truly influenced by the ego-vehicle. Fig. 4 presents how the observation length influenced model performance. Remarkably, in the PIE dataset, prediction accuracy initially improved and then declined as the observation time extended, whereas in the JAAD dataset, longer observation times consistently resulted in higher accuracy. This is because the PIE contains more diverse pedestrian samples from urban downtown areas, where behavior before crossing can change unpredictably. As a result, extended observation times may lead to misjudgments by the model. In contrast, pedestrian crossing behavior in JAAD is relatively simpler, allowing longer observation periods to more effectively capture pedestrian movement patterns.
Figure 3: Performance evaluations on the PIE (left) and JAAD (right) datasets across various TTE settings in both validation and test sets.
Figure 4: Impact of observation length on model performance across PIE (left) and JAAD (right) datasets.
We also tested the model’s performance with different reference line coordinates, as shown in Table II. Since the $(\Delta x_{t}-\Delta x_{t-1},\Delta y_{t}-\Delta y_{t-1})$ in PDM are only influenced by the line’s slope, we varied the vertical coordinates of points $A$ and $C$ for testing. With little variation in the line slope, the model’s performance remained less affected.
TABLE II: The impact of reference line coordinates.
@@ IV-C Evaluations
TABLE III: Ablation study on encoder components. Ours in bold.
@@
Acc 88.17/84.76 84.73/82.52 82.19/84.45 90.43/84.43 91.21/87.43
AUC 83.63/70.28 77.84/50.00 71.72/70.29 86.36/71.89 88.13/78.16
F1 75.04/52.41 66.78/0.00 57.98/52.17 79.57/54.16 81.61/64.00
Precision 75.08/57.71 68.94/0.00 65.81/56.42 80.53/55.82 80.97/64.11
TABLE IV: Impact on features used in $X_{pe}$. $P_{i}(C)$: coordinates variation in PDM, $P_{i}(R)$: area ratio in PDM, $P_{i}(\Delta d)$: depth variation. Best in bold and second-best underlined in each dataset.
@@
Acc 89.94/84.12 90.07/84.19 90.22/82.09 90.43/82.75 91.21/87.43 91.44/84.12
AUC 86.12/72.16 84.73/73.92 83.39/73.54 87.69/74.08 88.13/78.16 88.12/72.53
F1 78.80/54.22 78.08/56.25 77.35/54.11 80.33/55.19 81.61/64.00 81.93/54.65
Precision 78.76/54.66 81.92/54.50 85.81/49.00 78.30/50.57 80.97/64.11 82.01/54.58
TABLE V: Effect of motion features used in $X_{ev}$.
@@
90.22% 86.15% 79.18% 79.96% 78.42%
90.60% 85.52% 79.29% 83.04% 75.85%
91.21% 88.13% 81.61% 80.97% 82.26%
Improvement 0.99% 1.98% 2.43% 1.01% 3.84%
Figure 5: Visualizations of the predicted pedestrian crossing intentions. In the 4-row by 5-column grid, we show the observation frames ($T$) before the behavior in the first two rows and the frames at the start of the action in the last two rows, along with the correct predictions on the PIE and JAAD datasets. Failure cases are enclosed in dashed boxes, including: a) the ego-vehicle turning right (blue arrow) at a T-intersection while the pedestrian crosses vertically (white arrow); b) keypoint outliers; and c) the pedestrian waiting by the roadside.
To evaluate the effectiveness of each network component, ablation studies were conducted on the encoders and features used in $X_{pe}$ and $X_{ev}$, as shown in Table III - Table V. Left and right values in Table III and IV represent metrics in PIE and JAAD respectively. In Table III, the position encoder proved most effective in PIE, whereas skeletal information played a critical role in JAAD. The motion state of the ego-vehicle also significantly influenced pedestrian crossing behavior. In contrast to the results in Table I, incorporating skeletal information in PIE led to a 1-2% performance improvement. Overall, the model performs best when all encoders are combined. Table IV further illustrates the different features used in $X_{pe}$. Our proposed PDM brought a 4.26% F1 improvement in PIE and a 9.89% improvement in JAAD. Notably, regarding PIE, the PDM delivered an additional 1.28% F1 improvement when combined with the area ratio $R$. However, the inclusion of $P_{i}(C)$ in PDM improved less in JAAD compared to PIE, likely because the lateral movement can be more easily captured by other positional features. To compare performance with real depth estimation methods, we calculated the metric depth variation $\Delta d=d_{t}-d_{t-1}$ using Depth Anything [48] instead of $R$ in PDM, and achieved comparable results in PIE, while largely decreased in JAAD, due to the occlusion and errors in depth estimation. In Table V, we also demonstrated the effectiveness of acceleration in PIE, resulting in a 2.43% F1 improvement, with a particular increase in Recall.
Fig. 5 presents visualized predictions for the PIE and JAAD datasets, with the first two rows showing the observation period before TTE and the last rows depicting the action period. Additionally, failure cases are shown, typically falling into three categories: vertical crossing when the ego-vehicle is turning, keypoint outliers, and pedestrians waiting by the roadside. These areas can be addressed in future work.
V Conclusion
@@
In this study, we propose GTransPDM, a Graph-embedded Transformer with a Position Decoupling Module for pedestrian crossing intention prediction. It effectively fuses multi-modal features with position, ego-vehicle motion, and pedestrian pose, ensuring a lightweight, real-time application. The Position Decoupling Module captures positional variations with road-boundary-like reference lines and bounding box changes to decompose lateral movements and simulate depth in on-board image views. Additionally, a GCN-based Transformer models the spatial and temporal dynamics of pedestrian skeletons. Our method predicts pedestrian crossing behavior 1-2 seconds in advance, achieving 92% accuracy on the PIE dataset and 87% on the JAAD dataset, with a run time of 0.05ms, surpassing existing state-of-the-art solutions. For simplicity and efficiency, this work models road boundaries as constant using reference lines. Future work could improve this by incorporating semantic segmentation for dynamic road border representations.
ACKNOWLEDGMENT
This work was partly supported by the Scientific and Technological Developing Project of Jilin Province (20240402076GH); and the China Scholarship Council (202306170152). Antonio M. López acknowledges the financial support to his general research activities given by ICREA under the ICREA Academia Program. Antonio also thanks the Spanish grant PID2020-115734RB-C21 (MCIN/AEI/10.13039/501100011033) for the synergies generated for this paper. Antonio and Chen acknowledge the support of the Generalitat de Catalunya CERCA Program and its ACCIO agency to CVC’s general activities.

### v2

@@ III-A Datasets and Implementation Details
The effectiveness of the proposed method was evaluated using two large-scale datasets in the context of autonomous driving – JAAD [23] and PIE [24]. Both datasets are annotated with the crossing point in time, enabling the TTE calculation. The video frame rate from the two datasets is 30 FPS. Instead of ego-vehicle speed, five ego-vehicle motion states were used in the $\mathbf{X}_{ev}$ of JAAD, as velocity data are not available. The data used in the PIE experiments was split into train, validation (val), and test at a ratio of 0.5:0.1:0.4, with random splitting based on person IDs. For comparison with existing methods (Table I and II), the performance following the data split in [3] was also evaluated, with Set 01, 02, 04 for training, Set 05, 06 for validation and Set 03 for testing. Each video set contains a continuous driving recording segmented into multiple 10-minute clips. For the JAAD dataset, the JAAD_all subset and adhered to the default data split were used [3]. Furthermore, 0.5 second ($T=16$) observations were utilized to predict crossing behavior after 1-2 seconds ($TTE\in[30,60]$), sampled data based on the sliding window with an overlap of 0.6, 0.8 for PIE and JAAD dataset, respectively.
@@
In the network architecture, the dimension of three encoder layers $C_{d}$, hidden layers in GCN $d_{hid}$, and hidden size in Transformers are all set to 64. With the concatenate operation in the feature fusion block, the channel dimension of the fused feature is $1923\times C_{d}=$ 192. In the transformer encoder, $dropout=0.1$, 4 attention heads, and 4 stacked layers were applied to extract temporal information. The reference lines in PDM were set to one line between points: $A=(0,H)$ and $B=(W/2,Y_{min})$ while another between: $B=(W/2,Y_{min})$ and $C=(W,H)$, as shown in Fig. 2 (a), where $Y_{min}$ is the minimum $y_{t}$ of all samples in each dataset. The minor differences in coordinates have little impact on model performance in subsequent experiments, as shown in Table III. The keypoints were generated by the pre-trained Alphapose [25], while the additional neck, hip, and body center points were further calculated based on the average of adjacent points.
@@
Figure 3: Visualizations of predicted pedestrian crossing intentions in challenging scenarios. Displayed are the crossing probabilities with and without our proposed PDM, where values above 50% indicate crossing. Our method performs robustly under complex conditions, including irregular intersections (a) and curved roads (e), while also capturing long-range pedestrian behavior (b), and ambiguous roadside behavior (c) and (d). Failure cases typically falling into the parallel crossing behavior during vehicle turning, keypoints outliers, and the pedestrians yielding behavior to the ego-vehicle, as shown on the right.
@@
TABLE II: Model efficiency comparisons with recent models.
Model Acc (PIE) Acc (JAAD) Params (Millions) Inference Time (ms)
Pedestrian Graph+ [28] 0.89 0.86 0.07 5.47
PedGNN [32] 0.68 0.80 0.58
PCPA [3] 0.87 0.85 31.165 38.6
PIT [31] 0.91 0.87 21.2 4.8
GransPDM (w/o $X_{ke}$) 0.92 0.84 0.13 0.03
GransPDM 0.90 0.87 0.23 0.05
TABLE III: The impact of reference line coordinates.
@@
TABLE IV: Ablation study on multi-modalities. Ours in bold. Left: PIE dataset, Right: JAAD dataset.
@@
[…]
F1$\uparrow$ 75.04/52.41 66.78/0.00 57.98/52.17 79.57/54.16 81.61/64.00
[…]
TABLE V: Impact on features used in $X_{pe}$. $P_{i}(C)$: coordinates variation in PDM, $P_{i}(R)$: area ratio in PDM, $P_{i}(\Delta d)$: real depth variation. Best in bold and second-best underlined in each dataset.
@@
[…]
F1$\uparrow$ 78.80/54.22 77.35/54.11 80.33/55.19 81.61/64.00 81.93/54.65
[…]
III-B Evaluations
The effectiveness of our method was evaluated using Accuracy, Area Under Curve (AUC), F1 Score, Precision (P), and Recall (R), which are commonly used in existing studies. Table I and Table II compare our method and efficiency with recent approaches from the past five years, including Transformer-based IntFormer [10] and PIT [31], the graph-based Pedestrian Graph+[28], and several RNN-based methods. Except for Pedestrian Graph+ and BiPed, others show the same configurations as ours, following [3]. While PIT [31] captures traffic agent interactions using a ViT-based  [33] network with 16 stacked Transformer layers, incurring high computational costs, our lightweight network achieves superior performance, reaching 0.92 accuracy and 0.86 F1 in PIE using only position and ego-vehicle motion encoders, with 0.86 precision when incorporating the pose encoder. However, since [3] select samples following the sequence of long-period continuous videos, showcasing less randomness and performance degradation of $X_{ke}$ in our model. Thus, we conducted a 5-fold validation to further elucidate the contribution of $X_{ke}$ in Table VI of the Supplementary material and Table IV. Regarding JAAD, our method achieves an accuracy of 0.87 and a precision of 0.64, surpassing most existing approaches.
As an important component in the PDM, we evaluated the model’s performance with different reference line coordinates (see Table III). Since the values $(\Delta x_{t}-\Delta x_{t-1},\Delta y_{t}-\Delta y_{t-1})$ depend solely on the line’s slope, we varied the vertical coordinates of points $A$ and $C$ during testing. Our results indicate that slight variations in the line slope have minimal impact on performance, thereby underscoring the robustness in enhancing lateral pedestrian movement recognition. To evaluate the effectiveness of each network component, ablation studies were conducted on the modalities and features used in $X_{pe}$, as shown in Table IV and V. Left and right values represent metrics in PIE and JAAD respectively. In Table IV, the position encoder proves most effective in PIE, whereas skeletal information plays a critical role in JAAD. In contrast to the results in Table I, incorporating skeletal information in PIE leads to a 1-2% performance improvement. Overall, the model performs best when all encoders are combined. Table V further illustrates the different features used in $X_{pe}$. Our proposed PDM brings 4.26% and 9.89% F1 improvements in PIE and JAAD respectively. To compare it with real depth-based methods, we computed the metric depth variation $\Delta d=d_{t}-d_{t-1}$ using Depth Anything [34], a SOTA monocular depth estimation model, as a substitute for $R$ in PDM. While the model achieves comparable performance on the PIE dataset, it shows a significant drop on JAAD. This is mainly due to the limitations of monocular depth estimation, such as sensitivity to occlusion and failure in scenes with poor visibility. Fig. 3 presents challenging scenarios in PIE and JAAD datasets. Notably, in Fig. 3 (b), the ego-vehicle is traveling at high speed while the pedestrian remains at a distance, making it particularly challenging to detect the pedestrian’s approach. Owing to our PDM, the model reliably identifies crossing behavior, thereby ensuring a safe response by AVs. Beyond behavior prediction, the core idea of our model also facilitates human behavior understanding in mobile sensing scenarios, such as capturing fine-grained pose changes and providing a lightweight alternative to depth sensing.
IV Conclusion
@@
In this study, we propose GTransPDM, a Graph-embedded Transformer with a Position Decoupling Module (PDM) for pedestrian crossing intention prediction. It effectively fuses multi-modal features with position, ego-vehicle motion, and pedestrian pose, ensuring a lightweight, real-time application. The PDM enhances pedestrian lateral movement recognition and captures depth cues, while a GCN-enhanced Transformer models spatio-temporal skeleton dynamics. Our approach predicts crossing intent 1–2 seconds in advance, achieving 92% accuracy on PIE and 87% on JAAD, with a runtime of 0.05ms, outperforming existing state-of-the-art methods.